Annual starting salaries for college graduates with degrees in business administration are generally expected to be between $20,000 and $50,000. Assume that a 95% confidence interval estimate of the population mean annual starting salary is desired. Given the information in the Microsoft Excel Online file below, construct a spreadsheet to determine how large a sample should be taken for each desired margin of error.

For a margin of error of ± $300, the required sample size is n =

For a margin of error of ± $210, the required sample size is n =

For a margin of error of ± $120, the required sample size is n =
Answer #1

here from range rule of thumb ; estimated standard deviation =range/4=(50000-20000)/4=7500

For a margin of error of ± $300, the required sample size is n = (normsinv(0.975)*7500/300)^2=2401

For a margin of error of ± $210, the required sample size is n = (normsinv(0.975)*7500/210)^2=4900

For a margin of error of ± $120, the required sample size is n = (normsinv(0.975)*7500/120)^2=15006